### v1

# Not-so-glass-like Caging and Fluctuations of an Active Matter Model

Abstract
Simple active models of matter recapitulate complex biological phenomena. Their out-of-equilibrium nature, however, often makes these models beyond the reach of first-principle descriptions. This limitation is particularly perplexing when attempting to distinguish between different glass-forming mechanisms. We here consider a minimal active system in various spatial dimensions to identify the underlying processes at play. Activity is found to markedly impact cage escape processes and critical fluctuations, thus offering a distinct perspective on the role of activity sluggish dynamics.
Introduction –
Remarkably collective processes are observed in both macroorganisms, such as bird flocks and insect swarms Ramaswamy (2010); Marchetti et al. (2013); Cavagna and Giardina (2014), and microorganisms, such as bacterial swims and motile cells Elgeti et al. (2015); Polin et al. (2009); Ghosh et al. (2013); Maass et al. (2016). Equally remarkable is that the essence of these phenomena can be recapitulated by minimal models of active or self-propelled particles whose motion persists along a given direction Romanczuk et al. (2012); Bechinger et al. (2016); Zöttl and Stark (2016). Even these simplified models, however, remain challenging to describe from first principles. Hence, while the physical origin of motility-induced phase separation is by now fairly clear Cates and Tailleur (2015); Redner et al. (2013); Fily et al. (2014); Stenhammar et al. (2014); Digregorio et al. (2018); de Pirey et al. (2019), the nature of dynamical arrest at higher densities remains debated Ni et al. (2013); Berthier (2014); Nandi and Gov (2017); Liluashvili et al. (2017); Paul et al. (2023). The extent to which that arrest is akin to that of standard liquid glass formers, in particular, remains unclear.
Descriptions from either spin-glass theory Berthier and Kurchan (2013), mode-coupling theory (MCT) Janssen (2018); Farage and Brader (2014); Liluashvili et al. (2017); Szamel (2019); Reichert et al. (2021), or dynamical mean-field theory (DMFT) Agoritsas et al. (2019a, b); Morse et al. (2021) offer some physical insight into the matter, but also suffer from severe limitations. The first is based on a loose analogy, the second on uncontrolled approximations Pihlajamaa et al. (2023), and the third cannot be solved using existing numerical tools (other than in the dilute limit Manacorda and Zamponi (2022)). In that sense, the problem is even more challenging than that of simple glasses Charbonneau et al. (2017); Parisi et al. (2020). In this context, one promising approach is to employ numerical simulations to interpolate between systems of interest and the infinite-dimensional limit, $d\rightarrow\infty$, in which the mean-field description is exact. This strategy has been strikingly effective to elucidate the behavior of simple glass formers Parisi et al. (2020), notably by identifying the importance of marginality around the jamming transition Charbonneau et al. (2017), sorting out the algorithmic dependence of jamming Charbonneau and Morse (2023), and distinguishing between perturbative mean-field-like caging Biroli et al. (2022); Folena et al. (2022); Charbonneau et al. (2024); Laudicina et al. (2024) and non-perturbative instantons-like hopping Biroli et al. (2021a), around the dynamical arrest.
A particularly promising system for conducting such an analysis is an active Brownian particle random Lorentz gas (ABP-RLG) ten Hagen et al. (2011); de Macedo Biniossek et al. (2018); Zeitz et al. (2017); Sprenger et al. (2020); Janssen (2019); Zeitz et al. (2017). Recall that the RLG is a minimal model of transport in heterogeneous environments, in which a tracer evolves within spherical obstacles that are fixed and randomly distributed. Despite the apparent simplicity of this setup, even the standard RLG exhibits a rich phenomenology Höfling et al. (2006, 2008); Jin and Charbonneau (2015); Charbonneau et al. (2021); Biroli et al. (2021b, a). The RLG also has the same mean-field description as simple (hard-spheres-like) liquids in the limit $d\rightarrow\infty$ Biroli et al. (2021a); Manacorda and Zamponi (2022), and therefore distinguishes single-particle from genuinely collective phenomena in finite $d$. Prior studies of ABP-RLG have nevertheless mostly focused on the percolation regime Zeitz et al. (2017), which–given its purely geometrical nature–is unsurprisingly unchanged from that of the standard RLG. The glass-like arrest, by contrast, is expected to be much more sensitive to activity Ni et al. (2013); Berthier (2014); Liluashvili et al. (2017); Keta et al. (2022), but much less understood.
In this letter, with the help of simulations of the ABP-RLG in spatial dimensions $d=2\ldots 12$, we find that activity enhances the dynamical contribution of instanton-like cage escape processes and markedly alters particle-scale dynamical heterogeneity. In other words, activity changes the nature of dynamical arrest compared to standard glass formers.
Model and Methods –
Figure 1: Sample $d=2$ trajectory up to $t=30$ of a tracer with (a) $\widehat{\mathrm{Pe}}=0$ (BP-RLG) and (b) $10$ (ABP-RLG) within the same fixed obstacles at $\widehat{\varphi}=1.5$ (drawn to unit diameter). (c) Steady-state (bottom) radial distribution function of obstacles around the tracer, $g(r)$, and corresponding (top) effective potential for $\widehat{\mathrm{Pe}}=0,1,10$ (blue to red curves). The contact peak markedly grows with activity up to several $k_{\mathrm{B}}T$, thus reflecting an increase in the effective attraction between the tracer and obstacles.
The RLG model entails following the dynamics of a spherical tracer of radius $R$ within the void space left by fixed spherical obstacles of radius $R$, distributed uniformly at random at a (dimensionally-rescaled) volume fraction $\widehat{\varphi}=\rho R^{d}V_{d}/d$ for number density $\rho$, where $V_{d}$ is the volume of $d$-dimensional unit sphere Biroli et al. (2021a). (An equivalent system has a point tracer and obstacles of radius $2R$. In the following, $2R$ is set as the unit of length.) For the ABP-RLG, the tracer is assigned a constant intrinsic velocity $v_{0}$ along the direction of a unitary vector $\mathbf{e}(t)$ Howse et al. (2007). The associated overdamped Langevin equations with translational and rotational Gaussian white noise of zero mean and unit variance – $\mathbf{\xi}_{t}$ and $\mathbf{\xi}_{r}$, respectively –
$\langle\mathbf{\xi}_{r,t}(t)\rangle=0,\quad\langle\mathbf{\xi}_{r,t}(t)\cdot\mathbf{\xi}_{r,t}(t^{\prime})\rangle=\mathbf{\delta}(t-t^{\prime}),$ (1)
then have
$\dot{\mathbf{r}}(t)$ $=v_{0}\mathbf{e}(t)+\sqrt{2D_{t}}\mathbf{\xi}_{t}(t),$ (2)
$\dot{\mathbf{e}}(t)$ $=\sqrt{2D_{r}}\mathbf{\xi}_{r}(t)\otimes\mathbf{e}(t),$ (3)
where translational and rotational diffusion constants are related as $D_{r}=dD_{t}$, thus generalizing to arbitrary $d$ the natural choice made for $d=3$ systems Dhont (1996). This particular scaling, however, does not qualitatively alter the conclusions about the role of activity SI . Note that $\otimes$ corresponds to the cross product in $d=3$, and more generally denotes the projection of $\mathbf{\xi}_{r}(t)$ on the tangent surface perpendicular to $\mathbf{e}(t)$. Equation (3) therefore corresponds to $\mathbf{e}(t)$ moving on the surface of a unit hypersphere. Simulations leverage a standard event-driven Brownian dynamics scheme Foffi et al. (2005); Scala et al. (2007); Charbonneau et al. (2024) with the Brownian component giving rise to purely ballistic collisions, thus preventing tracer and obstacles from ever overlapping.
Activity strength is characterized by a Péclet number $\mathrm{Pe}$. Prior simulation works have defined it as the ratio of the diffusive time scale $4R^{2}/D_{t}$ and the ballistic time scale $2R/v_{0}$, i.e., $\mathrm{Pe}^{s}=\frac{2v_{0}R}{D_{t}}$ Zöttl and Stark (2016); Zeitz et al. (2017), whereas MCT studies have used $\mathrm{Pe}^{\mathrm{MCT}}=\frac{v_{0}^{2}}{D_{t}D_{r}}$ Liluashvili et al. (2017) to obtain an effective diffusion coefficient $D_{\mathrm{eff}}=D_{t}(1+\mathrm{Pe}^{\mathrm{MCT}})$. In order to relate the two, we here define
$\mathrm{Pe}=\frac{v_{0}}{\sqrt{D_{t}D_{r}}}$ (4)
and more specifically the dimensionally rescaled $\widehat{\mathrm{Pe}}=\mathrm{Pe}/\sqrt{d(d-1)}$. In all cases, $\widehat{\mathrm{Pe}}\ll 1$ corresponds to (translational) diffusion dominating – with a standard Brownian particle (BP) having $\widehat{\mathrm{Pe}}=0$ – while $\widehat{\mathrm{Pe}}\gg 1$ corresponds to active motion dominating. As illustrated in Fig. 1, a BP tracer samples void space uniformly with entropic bottlenecks appearing in narrow passages; an ABP tracer, by contrast, mostly explores the surface of obstacles and traverses narrow passages fairly easily, but tends to get trapped in wedges instead. The tracer can only leave a wedge if the overall velocity vector points outward or has at least one component parallel to the tangent surface, a process mostly controlled by the rotational component of the diffusion at high $\widehat{\mathrm{Pe}}$. (This effect becomes significant and intricate for $d\geq 3$, because determining the moving direction then requires considering multiple obstacles SI .) The marked increase in the effective attraction between the tracer and obstacles due to self-propulsion Berthier et al. (2019) can also be gleaned from the radial distribution function, which sharply peaks at contact as $\widehat{\mathrm{Pe}}$ increases (see Fig. 1(c)).
Results –
We first consider the evolution with time $t$ of the tracer mean squared displacement (MSD), $\Delta(t)=\langle|\mathbf{r}(t)-\mathbf{r}(0)|^{2}\rangle$. In free space Zöttl and Stark (2016),
$\Delta=2dD_{t}t+2v_{0}^{2}\tau_{r}t-2v_{0}^{2}\tau_{r}^{2}(1-e^{-\frac{t}{\tau_{r}}})$ (5)
where $\tau_{r}=\frac{1}{(d-1)D_{r}}$ is the persistence time of a spherical particle. With appropriate dimensional rescaling,
$\hat{\Delta}=\hat{t}+\widehat{\mathrm{Pe}}^{2}(\hat{t}-1+e^{-\hat{t}})$ (6)
where $\hat{t}=t/\tau_{r}$, and $\widehat{\Delta}=\frac{\Delta}{2dD_{t}\tau_{r}}$ is accordingly set so that a free-space BP tracer has $\widehat{\Delta}=\hat{t}$ SI .
The rescaled Péclet number then collapses the short- and long-time free-space diffusion regimes for all $d$,
$\widehat{\Delta}(\hat{t})=\begin{cases}\hat{t}+\frac{1}{2}\widehat{\mathrm{Pe}}^{2}\hat{t}^{2}&\mbox{for $\hat{t}\ll 1$}\\ (1+\widehat{\mathrm{Pe}}^{2})\hat{t}&\mbox{for $\hat{t}\gg 1$}.\end{cases}$ (7)
Put differently, the MSD of a free ABP is diffusive (tsimilar-$\sim t$) at very short times, ballistic ($\sim t^{2}$) at intermediate times, and diffusive with an enhanced effective diffusion coefficient at long times Bechinger et al. (2016).
Figure 2: (a) MSD for $\widehat{\mathrm{Pe}}=10$ for $\widehat{\varphi}=0.5,1,2,2.5,4$ (from top to bottom), and $d=3,4,6,12$ (fading as $d$ decreases) compared with the free-space result from Eq. (7) (black dashed lines). (Insets) The long-time diffusivity (plateau) in the ABP-RLG is enhanced relative to the BP-RLG at the same densities (from top to bottom) in $d=12$. (b) The critical density $\widehat{\varphi}_{\mathrm{d}}$ for $d=12$ (squares) increases with $\widehat{\mathrm{Pe}}$ and appears to saturate around the percolation threshold $\widehat{\varphi}_{\mathrm{p}}(d=12)=2.64(9)$ Biroli et al. (2021a) (black dashed line), and so does the scaling exponent $\gamma$ (circles). Results for $\widehat{\mathrm{Pe}}=0$ in the limit $d\rightarrow\infty$$\widehat{\varphi}_{\mathrm{d}}=2.4034$ and $\gamma=2.33786$ (empty symbols) – are provided as reference Biroli et al. (2021a); Kurchan et al. (2013). Lines are guides for the eye. (Inset) Critical scaling of the long-time diffusivity $\widehat{D}_{\infty}$ for $d=12$ with $\widehat{\mathrm{Pe}}=1,2,3,5,10$ (blue to red lines).
For the BP-RLG, the MSD cleanly converges to the $d\rightarrow\infty$ DMFT results as $d$ increases Charbonneau et al. (2024); Manacorda et al. (2020); for the ABP-RLG, the dimensional trend is similar, but DMTF predictions remain unavailable (see Fig. 2(a)). As expected, the long-time diffusion is enhanced as the Péclet number grows (see Fig. 2(a) inset). More surprisingly, this effect increases with density in the diffusive regime. In other words, when an abundance of obstacles geometrically hinders escapes, activity affects diffusion more by increasing the probability of exploring escape paths out of cages. Caging in active systems is therefore qualitatively different from that of standard liquids.
In order to estimate the resulting impact of this accelerated uncaging, we estimate the (avoided) dynamical transition $\widehat{\varphi}_{\mathrm{d}}$, at which the tracer would localize in an MCT-like description. The long-time diffusivity, $\widehat{D}_{\infty}=\lim\limits_{\hat{t}\rightarrow\infty}\frac{\widehat{\Delta}(\hat{t})}{\hat{t}}$, is therefore taken to scale as
$\widehat{D}_{\infty}\sim\left\lvert{\widehat{\varphi}-\widehat{\varphi}_{\mathrm{d}}}\right\rvert^{\gamma}.$ (8)
Because for $d\leq 8$, the geometric percolation transition takes place below the mean-field localization, i.e., $\widehat{\varphi}_{\mathrm{p}}\leq\widehat{\varphi}_{\mathrm{d}}$ Biroli et al. (2021a), the two phenomena are then easily confounded. We here sidestep this problem by considering the case $d=12$, where $\widehat{\varphi}_{\mathrm{p}}=2.64(9)$ is well above $\widehat{\varphi}_{\mathrm{d}}$ for the BP-RLG Biroli et al. (2021a). (The same holds for $d=10$ SI .) In such high $d$, perturbative deviations (in $1/d$) from mean-field $d\rightarrow\infty$ physics are also rather small Charbonneau et al. (2024). As expected from prior studies of active systems, the critical density grows with increasing Péclet number (Fig. 2(b)) Berthier (2014); Liluashvili et al. (2017). Interestingly, the effect eventually saturates at around 10%. This gap is much larger than the $\sim 1\%$ predicted by MCT for $d=2$ systems with comparable activity Liluashvili et al. (2017). The saturation further coincides with the percolation threshold, $\widehat{\varphi}_{\mathrm{p}}$, which provides a clear geometric upper bound to cage escapes. Put differently, by preferentially sampling low-dimensional features of caging, the active tracer sidesteps the entropic bottlenecks faced by the BP-RLG, and is hence largely oblivious to the associated instantons.
Figure 3: Time evolution of $\widehat{\alpha}_{2}$ of the ABP-RLG with (a) $\widehat{\mathrm{Pe}}=5$ and various obstacle densities $\hat{\varphi}=1,1.5,2,2.3,2.38,2.4,2.42$ (blue to red lines), and (b-g) $\widehat{\mathrm{Pe}}=10$ in $d=3,4,6,8,10,12$ (blue to red lines) for given obstacle densities. For $\widehat{\varphi}=0$, analytical expressions (dashed lines) can be obtained in all $d$ SI . Note that results for $\widehat{\varphi}\gtrsim\widehat{\varphi}_{\mathrm{p}}$, which do not diffuse at long times, are faded out to clarify the dimensional trend, and the ranges corresponding to percolation physics are dotted out.
The behavior of the exponent $\gamma$ is also physically interesting. Although its precise numerical value is coupled to that of $\hat{\varphi}_{d}$ and therefore depends on the details of the fitting range SI , it nevertheless robustly increases with $\widehat{\mathrm{Pe}}$. This growing sensitivity of the diffusion to activity is markedly distinct from its invariance assumed in the analysis of spin-glass models Berthier (2014), and opposite to recent MCT predictions Reichert et al. (2021). Whether these calculations considered the relevant length scale or properly accounted for the impact of activity on the structure is, however, unclear, so we cannot exclude that the discrepancy could be resolved through a more careful analysis of the theory. In any event, the ABP-RLG findings provide a stringent test of the impact of activity on sluggish dynamics.
In order to probe the nature of active dynamics more thoroughly, we also consider its dynamical fluctuations. In the RLG model, these are fully captured by the non-Gaussian parameter (NGP),
$\alpha_{2}(t)=\frac{d}{d+2}\frac{[\langle r^{4}(t)\rangle]}{[\langle r^{2}(t)\rangle]^{2}}-1,$ (9)
[…]
Figure 3 shows that at both short and long times $\widehat{\alpha}_{2}$ vanishes for the APB-RLG, as expected from the diffusive nature of transport in either regime. At intermediate times, the NGP of free particles turns negative due to the anti-correlation of displacement along different dimensions, as noted in Ref. ten Hagen et al. (2011). More interestingly, just like the BP-RLG, the active model exhibits a long-time peak as $\hat{\varphi}$ approaches $\hat{\varphi}_{d}$. Unlike that model, however, a separate short-time ($\hat{t}=\mathcal{O}(1)$) peak also develops and grows in that density regime. Keta et al. Keta et al. (2022) have suggested that the former corresponds to the universal behavior of liquids glass-formers while the latter reveals “the nontrivial influence of many-body interactions in persistent liquids”. The ABP-RLG results, however, clearly indicate that both effects are single-particle in nature.
The finite-time position of the first peak and its growth with dimension lead us to a direct geometrical interpretation. Figure 4 indeed reveals that the peak height first increases with $\hat{\varphi}$, saturates around $\hat{\varphi}_{d}$, and then decreases, and that $\widehat{\mathrm{Pe}}$ linearly enhances the effect. Given that $\hat{\alpha}_{2}$ encodes fluctuations in cage sizes, which are maximal around $\hat{\varphi}_{d}$ Biroli et al. (2022); Charbonneau et al. (2024), the first trend follows. It further highlights the critical origin of the phenomenon, with an actual divergence expected in the limit $d\rightarrow\infty$. The second trend is consistent with an active tracer exploring ever lower dimensional features – down to edges and wedges – of its surroundings upon approaching caging. The robustness and magnitude of this effect suggest its importance in capturing the dynamics of active glass formers. A proper theory of these systems would then be expected to capture this distinctive fluctuation-based signature.
Figure 4: Density evolution of the height of the first peak of $\hat{\alpha}_{2}$ at $\widehat{\mathrm{Pe}}=10$ for various $d$ and obstacle densities in the ABP-RLG. The peak height generally saturates as $d$ increases, but is roughly proportional to $d$ around $\hat{\varphi}_{d}$ as in BP-RLG Charbonneau et al. (2024). (inset) The linear growth of the peak height with $\widehat{\mathrm{Pe}}$ for $d=12$ and $\hat{\varphi}=1$ (dashed line), $2.5$ (solid line) and $4$ (dot-dashed line).
Conclusion –
Studying the ABP-RLG provides various key insights into the physics of active glasses. In particular, the mean-field–like dynamical transition shifts to higher densities with activity and saturates at the percolation transition, thus identifying activity as a mechanism to sidestep what would be instantonic (or activated) processes in standard glass formers. Activity also leads to the exploration of lower-dimensional caging features, thus resulting in a clear and distinct signature in the NGP around $\widehat{\varphi}_{\mathrm{d}}$. This finite-dimensional echo of mean-field criticality is well within the reach of simulation and experiments in these systems and is therefore a signature with which to contrast standard and active glass formers. Furthermore, these findings highlight the key importance of solving the DMFT, as well as its small and large fluctuations, as the necessary path toward achieving a first-principle description of glasses, whether active or not.
Acknowledgments–
We would like to thank Ludovic Berthier, Giampaolo Folena, Yi Hu, and Grzegorz Szamel for many stimulating exchanges. This work was supported by a grant from the Simons Foundation (Grant No. 454937). Data relevant to this work have been archived and can be accessed at the Duke Digital Repository https://doi.org/10.7924/XXXXXXX.