### v2

# Not-So-Glass-Like Caging and Fluctuations of an Active Matter Model

Abstract
Simple active models of matter recapitulate complex biological phenomena. The out-of-equilibrium nature of these models, however, often makes them beyond the reach of first-principle descriptions. This limitation is particularly perplexing when attempting to distinguish between different glass-forming mechanisms. We here consider a minimal active system in various spatial dimensions to identify the processes underlying their sluggish dynamics. Activity is found to markedly impact cage escape processes and critical fluctuations associated with exploring lower-dimensional caging features.
Introduction –
Remarkably collective processes are observed in both macroorganisms, such as bird flocks and insect swarms Ramaswamy (2010); Marchetti et al. (2013); Cavagna and Giardina (2014), and microorganisms, such as swimming bacteria and motile cells Elgeti et al. (2015); Polin et al. (2009); Ghosh et al. (2013); Maass et al. (2016). Equally remarkable is that the essence of these phenomena can be recapitulated by schematic models of active, or self-propelled, particles whose motion persists along a given direction Romanczuk et al. (2012); Bechinger et al. (2016); Zöttl and Stark (2016); Nandi et al. (2018); Mandal and Sollich (2020). Even these simplified models, however, remain challenging to describe from first principles. Hence, while the physical origin of motility-induced phase separation is now fairly clear Cates and Tailleur (2015); Redner et al. (2013); Fily et al. (2014); Stenhammar et al. (2014); Digregorio et al. (2018); de Pirey et al. (2019); Paoluzzi et al. (2022), the nature of dynamical arrest at higher densities remains debated Janssen (2018); Berthier et al. (2019); Janssen (2019). Active glasses are often deemed equivalent to their passive counterparts with quantitatively enhanced (effective) diffusion Howse et al. (2007); Farage and Brader (2014); Flenner et al. (2016) and temperature Berthier and Kurchan (2013); Nandi and Gov (2017); Cugliandolo et al. (2019), and longer ranged velocity correlations Szamel (2016); Caprini et al. (2020). Yet some of their phenomenological features, such as their dynamical heterogeneity Paul et al. (2023), are qualitatively distinct. The extent to which arrest in active systems is akin to that of standard liquid glass formers therefore remains unclear Berthier et al. (2017); Ni et al. (2013); Berthier (2014).
Descriptions from either spin-glass theory Berthier and Kurchan (2013), mode-coupling theory (MCT) Janssen (2018); Farage and Brader (2014); Liluashvili et al. (2017); Feng and Hou (2017); Szamel (2019); Reichert et al. (2021), or dynamical mean-field theory (DMFT) Agoritsas et al. (2019a, b); Morse et al. (2021) offer some insight into the matter, but also suffer from severe limitations. The first is based on a loose analogy, the second on uncontrolled approximations Pihlajamaa et al. (2023), and the third cannot be solved using existing numerical tools (other than in the dilute limit Manacorda and Zamponi (2022)). In that sense, the problem is even more challenging than that of simple glasses for which these descriptions have traditionally been used Charbonneau et al. (2017); Parisi et al. (2020). In this context, one promising approach is to employ numerical simulations to interpolate between systems of interest and the infinite-dimensional limit, $d\rightarrow\infty$, in which the mean-field description is exact. This analysis has been strikingly effective in elucidating the behavior of simple glass formers Parisi et al. (2020), notably by identifying the importance of marginality around the jamming transition Charbonneau et al. (2017), sorting out the algorithmic dependence of jamming Charbonneau and Morse (2023), and distinguishing between perturbative mean-field-like caging Biroli et al. (2022); Folena et al. (2022); Charbonneau et al. (2024); Laudicina et al. (2024) and nonperturbative instantonlike single-particle hopping Biroli et al. (2021a); Charbonneau et al. (2014); Dzero et al. (2009), around the dynamical arrest.
A particularly promising system for conducting such an analysis for active glass formers is the active Brownian particle random Lorentz gas (ABP-RLG) ten Hagen et al. (2011); de Macedo Biniossek et al. (2018); Zeitz et al. (2017); Sprenger et al. (2020); Janssen (2019); Zeitz et al. (2017). Like the passive RLG – a prototypical continuum-space model for kinetic theory Dorfman et al. (2021) and percolation Elam et al. (1984); Meester and Roy (1996); Rintoul and Torquato (1997); Charbonneau et al. (2021); Höfling et al. (2006, 2008); Jin and Charbonneau (2015); Charbonneau et al. (2021) – the ABP-RLG evolves a tracer within the void space left by fixed spherical obstacles. Interestingly, the resulting interplay between dynamics and geometry has the same mean-field description as simple (hard-sphere-like) liquids in the limit $d\rightarrow\infty$ Biroli et al. (2021a, b); Manacorda and Zamponi (2022), and can therefore be used to distinguish single-particle from genuinely collective phenomena in finite $d$. Prior studies of ABP-RLG have mostly focused on the percolation regime Zeitz et al. (2017), which–given its purely geometrical nature–is unsurprisingly unchanged from that of the standard RLG. The glasslike arrest, which is expected to be much more sensitive to activity Ni et al. (2013); Berthier (2014); Liluashvili et al. (2017); Keta et al. (2022), is much less understood.
In this Letter, using simulations of the ABP-RLG in spatial dimensions $d=2\ldots 12$, we find that activity results in lower-dimensional caging features being explored in the dynamically sluggish regime, thereby (i) enhancing escape processes by sidestepping instantonlike hopping, and (ii) giving rise to distinct critical fluctuations. In short, we find that activity alters the nature of the dynamical arrest from that of standard glass formers.
Model and methods –
Figure 1: Sample $d=2$ trajectory up to $t=30$ of a tracer with (a) $\widehat{\mathrm{Pe}}=0$ (BP-RLG) and (b) $10$ (ABP-RLG) within the same fixed obstacles at $\widehat{\varphi}=1.5$ (drawn to unit diameter). The passive tracer uniformly samples void space and rarely escapes the entropic bottleneck; the active tracer preferentially explores the surface of the obstacles and easily traverses the narrow path out. Activity alters the relaxation process. (c) Steady-state (bottom) radial distribution function of obstacles around the tracer, $g(r)$, and (top) corresponding effective potential for $\widehat{\mathrm{Pe}}=0,1,10$ (blue to red lines). The contact peak markedly grows with activity up to several $k_{\mathrm{B}}T$, indicating a stronger effective attraction between the tracer and the obstacles.
The RLG model follows the dynamics of a spherical tracer of radius $R$ within the void space left by fixed spherical obstacles also of radius $R$, distributed uniformly at random at a (dimensionally rescaled Biroli et al. (2021a, 2022); Zöttl and Stark (2016)) volume fraction $\widehat{\varphi}=\rho R^{d}V_{d}/d$ for number density $\rho$, where $V_{d}$ is the volume of $d$-dimensional unit sphere. For the ABP-RLG, the tracer is assigned a constant intrinsic velocity $v_{0}$ along the direction of a unitary vector $\mathbf{e}(t)$ Howse et al. (2007). The associated overdamped Langevin equations with translational $\mathbf{\xi}_{t}$ and rotational $\mathbf{\xi}_{r}$ Gaussian white noise of zero mean and unit variance,
$\langle\mathbf{\xi}_{r,t}(t)\rangle=0,\quad\langle\mathbf{\xi}_{r,t}(t)\cdot\mathbf{\xi}_{r,t}(t^{\prime})\rangle=\mathbf{\delta}(t-t^{\prime}),$ (1)
then have
$\dot{\mathbf{r}}(t)$ $=v_{0}\mathbf{e}(t)+\sqrt{2D_{t}}\mathbf{\xi}_{t}(t),$ (2)
$\dot{\mathbf{e}}(t)$ $=\sqrt{2D_{r}}\mathbf{\xi}_{r}(t)\otimes\mathbf{e}(t),$ (3)
where translational and rotational diffusion constants are related as $D_{r}=dD_{t}$, thus generalizing to arbitrary $d$ the natural choice made for $d=3$ systems Dhont (1996). Note that this particular choice of scaling, however, does not qualitatively alter the conclusions about the role of activity SI . Here $\otimes$ denotes the projection of $\mathbf{\xi}_{r}(t)$ on the tangent surface perpendicular to $\mathbf{e}(t)$. Equation (3) therefore corresponds to $\mathbf{e}(t)$ evolving on the surface of a unit hypersphere. Our simulations leverage a standard event-driven Brownian dynamics scheme Foffi et al. (2005); Scala et al. (2007); Charbonneau et al. (2024) with the Brownian component giving rise to purely ballistic collisions, thus preventing tracer and obstacles from ever overlapping.
Activity strength is characterized by a Péclet number $\mathrm{Pe}$. Prior simulations have defined this quantity as the ratio of the diffusive timescale $4R^{2}/D_{t}$ and the ballistic time scale $2R/v_{0}$, i.e., $\mathrm{Pe}^{s}=\frac{2v_{0}R}{D_{t}}$ Zöttl and Stark (2016); Zeitz et al. (2017), whereas MCT studies have used $\mathrm{Pe}^{\mathrm{MCT}}=\frac{v_{0}^{2}}{D_{t}D_{r}}$ Liluashvili et al. (2017) to obtain an effective diffusion coefficient $D_{\mathrm{eff}}=D_{t}(1+\mathrm{Pe}^{\mathrm{MCT}})$. In order to relate the two, we here empirically define
$\mathrm{Pe}=\frac{v_{0}}{\sqrt{D_{t}D_{r}}},$ (4)
and more specifically the dimensionally rescaled $\widehat{\mathrm{Pe}}=\mathrm{Pe}/\sqrt{d(d-1)}$, which recovers the expression for $D_{\textrm{eff}}$ in Eq. (7), and gives $\widehat{\mathrm{Pe}}\approx 1$ as crossover. That is, $\widehat{\mathrm{Pe}}\lesssim 1$ corresponds to (translational) diffusion dominating – with a standard Brownian particle (BP) having $\widehat{\mathrm{Pe}}=0$ – while $\widehat{\mathrm{Pe}}\gtrsim 1$ corresponds to active motion dominating SI . As illustrated in Fig. 1, a BP tracer samples void space uniformly with entropic bottlenecks appearing in narrow passages; an ABP tracer, by contrast, mostly explores the surface of obstacles and traverses narrow passages fairly easily, but can get wedged between obstacles, as previously noted in similar systems Bechinger et al. (2016). The tracer can only leave a cavity formed by obstacles if its velocity vector points outward or has at least one component parallel to the tangent surface, a process mostly controlled by the rotational component of the diffusion at high $\widehat{\mathrm{Pe}}$. (This effect becomes more significant and intricate for $d\geq 3$, because determining the moving direction then requires considering multiple obstacles at once SI .) The resulting increase in effective attraction between the tracer and the obstacles can be gleaned from the radial distribution function, which sharply peaks at contact as $\widehat{\mathrm{Pe}}$ increases (see Fig. 1(c)), consistent with observations from other active systems Berthier et al. (2019).
Results –
We first consider the evolution with time $t$ of the tracer mean squared displacement (MSD), $\Delta(t)=\langle|\mathbf{r}(t)-\mathbf{r}(0)|^{2}\rangle$. In free space Zöttl and Stark (2016),
$\Delta=2dD_{t}t+2v_{0}^{2}\tau_{r}t-2v_{0}^{2}\tau_{r}^{2}(1-e^{-\frac{t}{\tau_{r}}}),$ (5)
where $\tau_{r}=\frac{1}{(d-1)D_{r}}$ is the persistence time of a spherical particle. With appropriate dimensional rescaling,
[…]
where $\hat{t}=t/\tau_{r}$, and $\widehat{\Delta}=\frac{\Delta}{2dD_{t}\tau_{r}}$ is accordingly set so that a free-space BP tracer has $\widehat{\Delta}=\hat{t}$ SI . The rescaled Péclet number then collapses the short- and long-time free-space diffusion regimes for all $d$,
$\widehat{\Delta}(\hat{t})=\begin{cases}\hat{t}+\frac{1}{2}\widehat{\mathrm{Pe}}^{2}\hat{t}^{2}&\mbox{for $\hat{t}\ll 1$},\\ (1+\widehat{\mathrm{Pe}}^{2})\hat{t}&\mbox{for $\hat{t}\gg 1$}.\end{cases}$ (7)
Put differently, the MSD of a free ABP is diffusive ($\sim t$) at very short times, ballistic ($\sim t^{2}$) at intermediate times, and diffusive with an enhanced effective diffusion coefficient at long times Bechinger et al. (2016).
Figure 2: (a) MSD for $\widehat{\mathrm{Pe}}=10$ for $\widehat{\varphi}=0.5,1,2,2.5$, and 44 (from top to bottom), and $d=3,4,6$, and $12$ (fading as $d$ decreases) compared with the free-space result from Eq. (7) (black dashed line). (Insets) For $d=12$, (top) The ABP-RLG diffusivity at the same densities (lines from top to bottom) and (bottom) its ratio with the BP-RLG diffusivity. The long-time diffusivity of ABP-RLG is clearly enhanced. (b) The critical density $\widehat{\varphi}_{\mathrm{d}}$ for $d=12$ (squares) increases with $\widehat{\mathrm{Pe}}$ and appears to saturate around the percolation threshold $\widehat{\varphi}_{\mathrm{p}}(d=12)=2.64(9)$ Biroli et al. (2021a) (black dashed line), and so does the scaling exponent $\gamma$ (circles). Results for $\widehat{\mathrm{Pe}}=0$ in the limit $d\rightarrow\infty$$\widehat{\varphi}_{\mathrm{d}}=2.4034$ and $\gamma=2.33786$ (empty symbols) – are provided as reference Biroli et al. (2021a); Kurchan et al. (2013). Lines are guides for the eye. (Inset) Critical scaling of the long-time diffusivity $\widehat{D}_{\infty}$ for $d=12$ with $\widehat{\mathrm{Pe}}=1,2,3,5,10$ (blue to red lines).
For the BP-RLG, the MSD cleanly converges to the $d\rightarrow\infty$ DMFT results as $d$ increases Charbonneau et al. (2024); Manacorda et al. (2020); for the ABP-RLG, DMFT predictions remain unavailable, but the dimensional trend is similar (see Fig. 2(a)). As expected, the long-time diffusion is enhanced as the Péclet number grows (see Fig. 2(a) inset). More surprisingly, this effect increases with density in the diffusive regime. In other words, when an abundance of obstacles geometrically hinders escapes, activity affects diffusion more by increasing the probability of exploring escape paths out of cages. Caging in active systems is therefore qualitatively different than in standard liquids (see also Fig. 1).
In order to quantify the resulting impact of this accelerated uncaging, we estimate the (avoided) dynamical transition $\widehat{\varphi}_{\mathrm{d}}$, at which the tracer would localize in an MCT-like description. The long-time diffusivity, $\widehat{D}_{\infty}=\lim\limits_{\hat{t}\rightarrow\infty}\frac{\widehat{\Delta}(\hat{t})}{\hat{t}}$, is therefore taken to scale as
$\widehat{D}_{\infty}\sim\left\lvert{\widehat{\varphi}-\widehat{\varphi}_{\mathrm{d}}}\right\rvert^{\gamma}.$ (8)
Because for $d\leq 8$, the geometric percolation transition takes place below the mean-field localization, i.e., $\widehat{\varphi}_{\mathrm{p}}\leq\widehat{\varphi}_{\mathrm{d}}$ Biroli et al. (2021a), the two phenomena are then easily confounded. We here avoid this problem by considering the case $d=12$, where $\widehat{\varphi}_{\mathrm{p}}=2.64(9)$ is well above $\widehat{\varphi}_{\mathrm{d}}\approx 2.40$ for the BP-RLG Biroli et al. (2021a). (The same holds for $d=10$ SI .) In such high $d$, perturbative deviations (in $1/d$) from mean-field $d\rightarrow\infty$ physics are also rather small Charbonneau et al. (2024). As expected from prior studies of active systems, the critical density grows with increasing Péclet number (Fig. 2(b)) Berthier (2014); Liluashvili et al. (2017). Interestingly, the effect eventually saturates at around 10%. This gap is much larger than the $\sim 1\%$ predicted by MCT for $d=2$ systems with comparable activity Liluashvili et al. (2017). The saturation further coincides with the percolation threshold, $\widehat{\varphi}_{\mathrm{p}}$, which provides a clear geometric upper bound to cage escapes. By preferentially sampling low-dimensional features of caging, the active tracer effectively sidesteps the entropic bottlenecks of the BP-RLG in this regime Biroli et al. (2021a), and is therefore largely oblivious to these instantonlike hopping processes.
Figure 3: Time evolution of $\widehat{\alpha}_{2}$ for the ABP-RLG with (a) $\widehat{\mathrm{Pe}}=5$ and various obstacle densities $\hat{\varphi}=1,1.5,2,2.3,2.38,2.4,$ and $2.42$ (blue to red lines), and (b-g) $\widehat{\mathrm{Pe}}=10$ in $d=3,4,6,8,10,$ and $12$ (blue to red lines) for given obstacle densities. For $\widehat{\varphi}=0$, analytical expressions (dashed lines) can be obtained in all $d$ SI . Note that results for $\widehat{\varphi}\gtrsim\widehat{\varphi}_{\mathrm{p}}$, in which cases the traces does not diffuse at long times, are faded out to clarify the dimensional trend, and the ranges corresponding to percolation physics are dotted out.
The behavior of the exponent $\gamma$ is also physically interesting. Although its precise numerical value is coupled to that of $\hat{\varphi}_{d}$ and therefore depends on the details of the fitting procedure SI , it nevertheless robustly increases with $\widehat{\mathrm{Pe}}$. This growing sensitivity of the diffusion to activity is markedly distinct from its invariance assumed in the analysis of spin-glass models Berthier (2014), and opposite to recent MCT predictions Reichert et al. (2021). Whether these calculations considered the relevant length scale or properly accounted for the impact of activity on the structure is, however, unclear, so we cannot exclude that the discrepancy could be resolved through a more careful analysis of the theory. In any event, the ABP-RLG results provide a stringent test of the impact of activity on sluggish dynamics.
In order to probe the nature of active dynamics more thoroughly, we also consider its fluctuations. A common measure of particle-scale dynamical heterogeneity in standard glass-forming liquids Rahman (1964); Kob et al. (1997) is the non-Gaussian parameter (NGP), which in the RLG fully captures the effect Biroli et al. (2022); Folena et al. (2022),
$\alpha_{2}(t)=\frac{d}{d+2}\frac{[\langle r^{4}(t)\rangle]}{[\langle r^{2}(t)\rangle]^{2}}-1.$ (9)
Here, $\langle\cdots\rangle$ denotes averaging over initial conditions in one realization of disorder and $\cdots$] denotes averaging over multiple realizations. The quantity is dimensionally rescaled as $\widehat{\alpha}_{2}=d\cdot\alpha_{2}$ Charbonneau et al. (2024), thus capturing the small (perturbative) deviations from mean-field $d\rightarrow\infty$ physics Charbonneau et al. (2024). Deviations from $\widehat{\alpha}_{2}(t)=0$, which originate either from an enhanced displacement along a given direction or from directional correlations Biroli et al. (2022), are positive for a (van HoveHansen and McDonald (2006)) distribution of tracer displacements at fixed time with a large tail and negative otherwise. In passive systems, directional correlations result in the NGP markedly growing and peaking at times that grow upon approaching $\hat{\varphi}_{d}$. In active systems, its behavior is markedly different.
Figure 3 shows that at both short and long times $\widehat{\alpha}_{2}$ vanishes for the ABP-RLG, as expected from the diffusive nature of transport in either regime. At intermediate times, the NGP of free particles turns negative due to the anti-correlation of displacement along different dimensions, as previously noted in Ref. ten Hagen et al. (2011). Interestingly, just like the BP-RLG, the ABP-RLG exhibits a long-time peak which diverges as $\hat{\varphi}$ approaches $\hat{\varphi}_{d}$, as noted in Ref. Keta et al. (2022), but overlooked until then ten Hagen et al. (2011); Zheng et al. (2013); Ding et al. (2017). Even more interestingly, a separate short-time ($\hat{t}=\mathcal{O}(1)$) peak also develops and grows in that density regime. Keta et al. Keta et al. (2022) have suggested that the former peak corresponds to the universal behavior of liquid glass formers while the latter reveals “the nontrivial influence of many-body interactions in persistent liquids.” The ABP-RLG results, however, clearly indicate that both effects are single particle in nature.
The finite-time position of the first peak and its growth with dimension hint at a direct geometrical interpretation. Figure 4 reveals that the peak height first increases with $\hat{\varphi}$, saturates around $\hat{\varphi}_{d}$, and then decreases. It also shows that $\widehat{\mathrm{Pe}}$ linearly enhances the effect (inset) SI . Because $\hat{\alpha}_{2}$ encodes fluctuations in cage sizes, which are maximal around $\hat{\varphi}_{d}$ Biroli et al. (2022); Charbonneau et al. (2024), the explanation for the peak position naturally follows. It further highlights the critical origin of the phenomenon, with an actual divergence expected in the limit $d\rightarrow\infty$. The enhancement of the effect with $\widehat{\mathrm{Pe}}$ is consistent with an active tracer exploring ever lower dimensional features – down to edges and cavities formed by obstacles – of its surroundings upon approaching caging. The robustness and magnitude of this effect suggest its importance in capturing the dynamics of active glass formers. A proper theory of these systems would then be expected to capture this distinctive fluctuation-based signature.
Figure 4: Density evolution of the height of the first peak of $\hat{\alpha}_{2}$ for the ABP-RLG at $\widehat{\mathrm{Pe}}=10$ for various $d$ and obstacle densities. The peak height generally saturates as $d$ increases, but is roughly proportional to $d$ around $\hat{\varphi}_{d}$ similarly to the long-time peak of the BP-RLG Charbonneau et al. (2024). (inset) The linear growth of the peak height with $\widehat{\mathrm{Pe}}$ for $d=12$ and $\hat{\varphi}=1$ (dashed line), $2.5$ (solid line) and 44 (dot-dashed line).
Conclusion –
The ABP-RLG provides key insights into the physics of dense active systems. First, its mean-field–like dynamical transition shifts to higher densities with activity and saturates at the percolation transition, thus identifying activity as a mechanism that sidesteps the instantonlike (or activated) hopping processes of standard glass formers. Second, an active tracer tends to explore lower-dimensional caging features, resulting in a clear and distinct short-time peak of the NGP around $\widehat{\varphi}_{\mathrm{d}}$. This finite-dimensional echo of mean-field criticality is well within the reach of simulation and experiments in these systems and is therefore a distinctive feature with which to contrast standard and active glass formers. These findings could further be used to design systems in which the efficacy of active particle caging can be tuned, especially for $d>2$. Finally, these findings once again highlight the importance of solving the DMFT, as well as its small and large fluctuation corrections, as the necessary path toward achieving a first-principle description of glasses, both active and not.
Acknowledgements.
Acknowledgments–
We thank Ludovic Berthier, Giampaolo Folena, Yi Hu, Grzegorz Szamel, and Yikang Zhang for many stimulating exchanges. This work was supported by a grant from the Simons Foundation (Grant No. 454937). D. K. thanks the European Union Horizon 2020 research and innovation program under Grant Agreement No. 857470 and from the European Regional Development Fund via the Foundation for Polish Science International Research Agenda PLUS program GrantNo. MAB PLUS/2018/8, and the Ministry of Science and Higher Education’s initiative “Support for the Activities of Centers of Excellence Established in Poland under the Horizon 2020 Program” (agreement no. MEiN/2023/DIR/3795
Data availability –
Data relevant to this work have been archived and can be accessed at the Duke Research Data Repository Zheng et al. (2025).